OpenStudy (anonymous):

A physician has ordered 0.50 mg of atropine, intramuscularly. If atropine were available as 0.25 mg/mL of solution, how many milliliters would you need to give?

OpenStudy (anonymous):

I got it no worries

OpenStudy (aaronq):

For this question you need to use the definition of density: \(\sf density=\dfrac{mass}{volume}\) You're given a mass (in mg) and a density (in mg/mL) and you are asked to find the volume (in mL) that corresponds to these values, so you rearrange the formula: \(\sf density=\dfrac{mass}{volume}\rightarrow volume=\dfrac{mass}{density}\)

OpenStudy (aaronq):

lol aight